# InkSight: Offline-to-Online Handwriting Conversion by Learning to Read and Write

### v2

@@ 4.6 Ablation Studies
Recognition tasks improve derendering quality (Row 6-8).
Removing recognition tasks from the multi-task training mixture, while maintaining the relative ratios of other tasks, notably reduces overall performance across all evaluation datasets, particularly impacting accuracy which reflects the model’s semantic understanding ability.
@@
Unfreezing the Vision Transformer (ViT) in multi-task training typically incurs training instability, as evidenced by significant variance in model evaluations and an increased tendency to misinterpret background noise as textual content (see Appendix D). The observed increase in F1 scores can be attributed to the model interpreting more background noise as strokes, which leads to digital inks that closely mimic the style of the input images, resulting in a higher likelihood of overlapping bounding boxes in F1 calculations. More details in Section D.1.
5 Conclusion and Future Work
@@
In this work we present the first approach to converting photos of handwriting into digital ink. We propose a training setup that works without paired training data, which can be difficult to acquire. We show that our method is robust to a variety of input conditions, can work on full handwritten notes, and generalizes to out-of-domain sketches to some extent. Furthermore, our approach does not require complex modelling and can be constructed from standard building blocks. Future work may address the main limitation of our model which is the need to have a page segmentation to run derendering on the individual words identified by the segmentation.
6 Impact Statements
6.1 Ethical Considerations
Proposed model can not be used for open-ended generation.
We verify that proposed models are unable to perform open-ended ink generation. We provide random texts in task prompts of task Derender with Text (in Figure 3) with input images either empty or contain handwritings that do not match the label. We observe that the generated inks do not match the labels for 100 cases inspected, signifying that the model is unable to generate the prompted text that is not present in the input image. More details in Section L.2.
6.2 Future Societal Consequences
Our work could allow access to the digital ink underlying the physical notes, potentially enabling the training of better online handwriting recognizers for languages that are historically low-resource in the digital ink domain.
7 Acknowledgement
@@
The authors thank Leandro Kieliger for the help on synthetic rendered ink rendering, feedback on initial versions of the paper, Philippe Schlattner on human evaluation protocol building, Anastasia Fadeeva, Mircea Trăichioiu for the multiple discussions on weekly meetings, Efi Kokiopoulou, Diego Antognini, Henry Rowley, Reeve Ingle, Manuel Drazyk for feedback on initial versions of the paper, Sebastian Goodman, Jialin Wu for the help on implementing deterministic training and PaLI related questions, Xiao Wang on ViT related questions, Tom Duerig and Tomáš Ižo for leadership support.
Appendix A Full-page Results
@@
We show the full-page results on samples that resemble real-life inputs (mostly from Unsplash with keyword search for “handwriting”, and others collected by the authors with consent from the writers) produced by three variants of our models.
@@
Appendix B Data statistics and Preprocessing
We present the language distribution of the digital ink datasets used to train our in-house (-i suffix) and publicly available (-p suffix) models in Figures. 33, 34 and specific datasets and corresponding number of samples used in training our public model in Table 5
@@ Appendix B Data statistics and Preprocessing
Task Type Dataset Number of Samples
Derendering DeepWriting (words) 89,565
DeepWriting (lines) 33,933
DeepWriting (characters) 359,643
VNOnDB 66,991
SCUT-COUCH Chinese characters 1,998,784
SCUT-COUCH Chinese pinyin 156,53
OCR IAM word-level (train) 53,839
IMGUR5k (train) 181,792
RIMES word-level (train) 51,738
HierText (train) 5,978
ICDAR-2015 (train) 1,535
Appendix C Survey of Derendering Models
@@
Work Has code Has data Has model weights Year
@@
Chen et al. (2022b)
repository missing core elements:
data preprocessing, outdated/incompetable config files
2022
@@
Bhunia et al. (2021)
repository missing core elements:
training scripts, data preprocessing
2021
@@
Appendix D Further Details of Ablation Studies
D.1 Frozen ViT helps training stability
We investigated the impact of freezing the vision encoder during training on model stability and derendering performance. To assess training stability, we analyzed the ratio of empty ink predictions across training steps on the “golden” human traced dataset (detailed in Section 4.5.1). Empty inks typically occur when the model confuses tasks, mistakenly outputting text instead of ink or both (different tasks depicted in Figure 3). Our analysis revealed significantly greater variance between runs with an unfrozen setup compared to the consistent learning observed with a frozen vision encoder (Figure 35).
@@ D.1 Frozen ViT helps training stability
D.2 Inference type matters
As we first illustrated in Section 4.6, different inference tasks can produce different output digital inks especially when the input image of text is semantically or geometrically ambiguous due to either image quality or how the texts were written. For example the first column of Figure 37 where the word ”wich” is removed by the writer with a strike through line but still recognized by both Recognize and Derender (model intrinsic) and Derender with Text (extrinsic OCR system) inference, and the difference in understanding and locating the textual information in input image results in different output digital inks. Additionally, the output from Vanilla Derender where semantics are not intentionally involved during training is more robust to ambiguous inputs but shows poorer semantic consistency with the input image.
@@ D.2 Inference type matters
Appendix E Rendering and Data Augmentation
To create synthetic rendered inks of our online samples, we use the Cairo graphics library to render online ink samples onto a $224\times 224$ canvas. We then add several augmentations to these rendered samples, to ensure that they are closer in domain to the real-world samples expected.
@@ Appendix E Rendering and Data Augmentation
Rotation
angle (rad): $[-\frac{\pi}{4},\frac{\pi}{4}]$
Stroke color
RGB: $[0,1]^{3}$
Background color
RGB: $[0,1]^{3}$
Stroke Width
width: $[1\text{px},12\text{px}]$
Lines
line width: $[1\text{px},6\text{px}]$ line dist: $[10\text{px},100\text{px}]$ line color: RGB: $[0,1]^{3}$
Grids
line width: $[1\text{px},6\text{px}]$ line dist: $[10\text{px},100\text{px}]$ line color: RGB: $[0,1]^{3}$
Gaussian Noise
standard dev.: $[50,500]$
Box blur
radius: $[0\text{px},5\text{px}]$
@@
Figure 39: Demonstration of possible ways to combine all used augmentations. (Sample from Aksan et al. 2018)
Appendix F Additional Model Details
@@
Large-i ViT-L 303M
In-house proprietary datasets
mT5-Large 783M
as shown in Appendix B.
Small-i ViT-B 85M
In-house proprietary datasets
mT5-Base 247M
as shown in Appendix B.
Small-p ViT-B 85M
Publicly available datasets
mT5-Base 247M
as shown in Appendix B.
@@
Similar to PaLI models (Chen et al., 2022a, 2023a, 2023b), our models together with the training mixtures are implemented in JAX/Flax (Bradbury et al., 2018) using the open-source T5X, SeqIO (Roberts et al., 2023) and Flaxformer (Heek et al., 2023) frameworks. For training, we use the Adafactor optimizer (Shazeer & Stern, 2018) with $\beta_{1}=0$, second order moment of $0.8$, and a language-model–style teacher forcing with softmax cross-entropy loss. For the learning rate schedule, we use the linear decay scheduler with a peak learning rate of $0.01$, 5k steps of warmup and a linear decay with decay factor of $63e-$ 6, and a dropout of $0.1$ on both ViT encoder and the mT5 encoder-decoder. We train our models for 340k steps with batch size $512$. With frozen ViT encoders, the training of Small-i takes $\sim$33h on 64 TPU v5e chips and the training of Large-i takes $\sim$105h on 64 TPU v5e chips.
@@
Appendix G General Virtual Sketching Framework Performance
@@
Appendix H Online Handwriting Recognizer Details
@@
Parameter IAMOnDB
Small-i
IAM
derendered
Small-i
IAMOnDB +
IAM
derendered
Small-p
IAM
derendered
Small-p
IAMOnDB +
IAM
derendered
Large-i
IAM
derendered
Large-i
IAMOnDB +
IAM
derendered
Dropout 0.25 0.25 0.3 0.3 0.25 0.25 0.25
Learning rate 0.005 0.001 0.001 0.005 0.005 0.001 0.001
Appendix I Limitations
While our system exhibits satisfactory performance across various tested inputs, we have identified specific failure patterns that warrant further attention. As some examples shown in the Figures 38 and 36 we notice model’s performance deteriorates when processing samples with wider stroke widths or significant variations in stroke width. Additionally, we observed that the model’s ability to preserve handwriting details improves with the scaling of the vision component (visualized in Figure 5, aligned with findings in Section 4.5.1). However, we also noticed a lack of proportional scaling between the vision (ViT) and text-ink (mT5) components negatively impacts semantic understanding.
Furthermore, our investigation, as detailed in Section 4.6, demonstrated that unfreezing the vision heads allows the model to capture more input image details. Nevertheless, this enhancement comes at the cost of training instability and increased confusion between textual information and background noise. Future research in this domain may explore fine-tuning the vision backbone while leveraging the existing semantic understanding of ink representations, following a similar idea as described in (Zhai et al., 2022).
Appendix J Human Evaluation
@@ Appendix I Limitations
Task setup
You will be looking at input images of photos of handwriting in a variety of styles and the corresponding digital ink tracing performed either by a human or by one of our derendering models.
Each stroke is rendered in a different color and with color gradient (darker $\to$ whiter) to reflect the direction.
You will see the derendered ink as well as the derendered ink overlaid on top of the image and can use either to make the judgment.
For each sample answer the two questions on the right and click submit to proceed to the next sample.
Is Reasonable
Is the digital ink a reasonable tracing of the input image?
  • Yes, it’s a good tracing
  • It’s an okay tracing, but has some small errors
  • It’s a bad tracing, has some major artifacts
One interpretation of the quality of the ink could be like this:
If you were to use it in a note taking app, would you
  • keep it as is: good tracing
  • need to make some edits: okay tracing
  • not use it at all: bad tracing
If you need some examples to get a better feeling, you can use the ones below, however there is some ambiguity within the task since we don’t want to impose any strict rules but allow for a natural variety of what people find a good derendering (it is designed to be subjective).
Is Real
Could this digital ink have been produced by a human?
  • Yes
  • No
In this question go with your intuition, do you think the ink could have been traced by a [reasonable] human?
@@
Appendix K Model Card
@@
Model Architecture
A multimodal sequence-to-sequence Transformer (Vaswani et al., 2017) model with the mT5 (Xue et al., 2021) encoder-decoder architecture. It takes text tokens and ViT (Dosovitskiy et al., 2021) dense image embeddings as inputs to an encoder and autoregressively predicts discrete text and ink tokens with a decoder.
Input(s)
A pair of image and text.
Output(s)
Generated digital ink.
@@
Application
The model is for research prototype, and the public version is planned to be released and available for the public.
Known Caveats
None.
@@
System Description
This is a standalone model.
Upstream Dependencies
None.
Downstream Dependencies
None.
@@
Hardware & Software
Hardware: TPU v5e.
@@
For implementation details please refer to Appendix F.
Compute Requirements
Please refer to Appendix F.
@@
Training Datasets
The ViT encoder (Dosovitskiy et al., 2021) is pretrained on ImageNet-21k (Deng et al., 2009), mT5 encoder and decoder are initialized from scratch. The entire model is trained on the mixture of publicly available datasets described in Appendix B.
@@
Evaluation Methods
Human evaluation (reported in Section 4.5.1) and automated evaluations (reported in Section 4.5.2).
@@
Sensitive Use
The model is capable of converting images to digital inks. This model should not be used for any of the privacy intruding use cases, e.g. forging handwritings.
Known Limitations
Reported in Appendix I.
Ethical Considerations & Potential Societal Consequences
Reported in Sections 6.1 and 6.2.
Appendix L Additional Visualizations
@@
L.1 More Visualizations on Public Evaluation Datasets
@@
L.2 Open-ended Generation

### v4

@@ 4.6 Ablation Studies
Recognition tasks improve derendering quality (Rows 6-8).
Removing recognition tasks from the multi-task training mixture notably reduces overall performance across all evaluation datasets, particularly impacting accuracy which reflects the model’s semantic understanding ability. We hypothesize that the increase in F1 and drop in accuracy (such as the one observed on IMGUR5K) may be attributed to the model focusing less on the textual content and interpreting more background noise as strokes, which leads to digital inks that closely mimic the style of the input images, resulting in a higher likelihood of overlapping bounding boxes in F1 calculations.
@@
Unfreezing the Vision Transformer (ViT) in multi-task training typically incurs training instability, as evidenced by increase in variance in model evaluations and an increased tendency to misinterpret background noise as textual content. Explaining the observed increase in F1 scores in Section 4.6 can be attributed to the model misinterpreting background noise as strokes, generating digital ink that visually closely resemble the input images, including the noise. This behavior results in a higher likelihood of overlapping bounding boxes, inflating F1 calculations. More details in Section E.1.
4.7 Limitations
We acknowledge the following two main limitations of our approach:
  1. (1)
    Due to the limits in context length of transformer-based models, it can derender limited amount of characters or words in one model inference call; While our system incorporates efficient design choices such as resampling and sequence reduction to support full-page derendering, this remains a constraint.
  2. (2)
    Consequently, it requires some form of segmentation of the handwritten page to run derendering on rather short individual segments (words in our implementation) and in the current form cannot work on complex sketches or other objects with lots of strokes, extending it to handle complex sketches or dense stroke-based objects presents an exciting direction for future work.
One thing that is not a limitation of our model is the use of OCR to provide the word labels as shown in the first 2 rows of Table 4.6, the model performance degrades only slightly if the label is not provided or is inferred by the model. We also discuss some specific failure cases in Appendix J.
5 Conclusion
@@
In this work we present the first approach to converting photos of handwriting into digital ink. We propose a training setup that works without paired training data, which can be difficult to acquire. We show that our method is robust to a variety of inputs, can work on full handwritten notes, and generalizes to out-of-domain sketches to some extent. Furthermore, our approach does not require complex modeling and can be constructed from standard building blocks. We release the weights and inference code of the Small-p model, with a subset of model-generated and human expert-traced digital ink data. Future work may address the main limitations of our model stated above.
Acknowledgments
@@
The authors thank Leandro Kieliger for the help on synthetic ink rendering, feedback on initial versions of the paper, Philippe Schlattner on human evaluation protocol, Anastasiia Fadeeva, Mircea Trăichioiu for the multiple discussions, Efi Kokiopoulou, Diego Antognini, Henry Rowley, Reeve Ingle, Manuel Drazyk for feedback on initial versions of the paper, Sebastian Goodman, Jialin Wu for the help on implementing deterministic training and PaLI related questions, Xiao Wang on ViT related questions, Tom Duerig and Tomáš Ižo for leadership support.
Appendix A Appendix
B   Full-page Results B
C   Data Statistics and Preprocessing C
D   Survey of Derendering Models D
E   Further Details of Ablation Studies E
F   Rendering and Data Augmentation F
G   Additional Model Details G
H   General Virtual Sketching Framework Performance H
I   Online Handwriting Recognizer Details I
J   Expansion on Failure Cases J
K   Expansion on Out-of-Domain Generalization K
L   Human Evaluation L
M   Model Card M
N   Additional Visualizations N
O   Impact Statements O
supporting the main findings and methodology presented in this work.
Appendix B Full-page Results
@@
We show the full-page results on samples that resemble real-life inputs (mostly from Unsplash with keyword search for “handwriting", and others collected by the authors with consent from the writers) produced by three variants of our models.
@@
Figure 38: Large-i full page results of handwritten notes in a real-life scenario in Korean, credit: Unsplash.
Figure 39: Small-i full page results of handwritten notes in a real-life scenario in Korean, credit: Unsplash.
Figure 40: Small-p full page results of handwritten notes in a real-life scenario in Korean, credit: Unsplash.
Appendix C Data statistics and Preprocessing
We present the language distribution of the digital ink datasets used to train our in-house (-i suffix) and publicly available (-p suffix) models in Figures 41 and 42, datasets and corresponding number of samples used in training our public model in Table 4. The in-house OCR dataset comprises of approximately 500k samples, and in-house Digital Ink dataset comprises of approximately 16M samples.
@@ Appendix C Data statistics and Preprocessing
Task Type Dataset Samples License
Derendering DeepWriting (words) 89,565 Custom11footnotemark: 1
DeepWriting (lines) 33,933 Custom11footnotemark: 1
DeepWriting (characters) 359,643 Custom11footnotemark: 1
VNOnDB 66,991 MIT22footnotemark: 2
SCUT-COUCH (chars) 1,998,784 Custom33footnotemark: 3
SCUT-COUCH (pinyin) 15,653 Custom33footnotemark: 3
Recognition IAM word-level (train) 53,839 Custom44footnotemark: 4
IMGUR5k (train) 181,792 CC BY-NC55footnotemark: 5
RIMES word-level (train) 51,738 CC BY-NC55footnotemark: 5
HierText (train) 5,978 CC BY-SA66footnotemark: 6
ICDAR-2015 (train) 1,535 Custom77footnotemark: 7
  • 1
    Custom license (CC BY-NC-SA 4.0 + non-distribution). See: URL.
  • 2
    MIT License. See: URL.
  • 3
    Custom license (research-only, attribution). See: URL.
  • 4
    Custom license (research-only, non-commercial, attribution). See: URL.
  • 5
    CC BY-NC 4.0. See: URL.
  • 6
    CC BY-SA 4.0. See: URL.
  • 7
    Custom license (research-only, attribution). See: URL.
Appendix D Survey of Derendering Models
@@
Study Code Availability Data Availability Model Weights Availability Year
@@
Chen et al. (2022b) Repo missing core elements: data preprocessing, incompatible configs 2022
@@
Bhunia et al. (2021) Repo missing core elements: training scripts, data preprocessing 2021
@@
Appendix E Further Details of Ablation Studies
E.1 Frozen ViT helps training stability
We investigated the impact of freezing the vision encoder during training on model stability and derendering performance. To assess training stability, we analyzed the ratio of empty ink predictions across training steps on the “golden" human traced dataset (detailed in Section 4.5.1). Empty inks typically occur when the model confuses tasks, mistakenly outputting text instead of ink or both (different tasks depicted in Figure 3). Our analysis revealed significantly greater variance between runs with an unfrozen setup compared to the consistent learning observed with a frozen vision encoder (Figure 43).
@@ E.1 Frozen ViT helps training stability
E.2 Inference type matters
As we first illustrated in Section 4.6, different inference tasks can produce different output digital inks especially when the input image of text is semantically or geometrically ambiguous due to either image quality or how the texts were written. For example the first column of Figure 45 where the word "wich" is removed by the writer with a strike through line but still recognized by both Recognize and Derender (model intrinsic) and Derender with Text (extrinsic OCR system) inference, and the difference in understanding and locating the textual information in input image results in different output digital inks. Additionally, the output from Vanilla Derender where semantics are not intentionally involved during training is more robust to ambiguous inputs but shows poorer semantic consistency with the input image.
@@ E.2 Inference type matters
Appendix F Rendering and Data Augmentation
To create synthetic rendered inks of our online samples, we use the Cairo graphics library to render online ink samples onto a $224\times 224$ canvas. We then add several augmentations to rendered samples.
@@ Appendix F Rendering and Data Augmentation
Rotation angle (rad): $[-\frac{\pi}{4},\frac{\pi}{4}]$ [Uncaptioned image]
Stroke color RGB: $[0,1]^{3}$ [Uncaptioned image]
Background color RGB: $[0,1]^{3}$ [Uncaptioned image]
Stroke Width width: $[1\text{px},12\text{px}]$ [Uncaptioned image]
Lines line width: $[1\text{px},6\text{px}]$ line dist: $[10\text{px},100\text{px}]$ line color: RGB: $[0,1]^{3}$ [Uncaptioned image]
Grids line width: $[1\text{px},6\text{px}]$ line dist: $[10\text{px},100\text{px}]$ line color: RGB: $[0,1]^{3}$ [Uncaptioned image]
Gaussian Noise standard dev.: $[50,500]$ [Uncaptioned image]
Box blur radius: $[0\text{px},5\text{px}]$ [Uncaptioned image]
@@
Figure 47: Demonstration of possible ways to combine all used augmentations. (Sample from (Aksan et al., 2018))
Appendix G Additional Model Details
@@
Large-i ViT-L 303M In-house proprietary datasets
mT5-Large 783M as shown in Appendix C.
Small-i ViT-B 85M In-house proprietary datasets
mT5-Base 247M as shown in Appendix C.
Small-p ViT-B 85M Publicly available datasets
mT5-Base 247M as shown in Appendix C.
@@
Similar to PaLI models (Chen et al., 2022a; 2023a; 2023b), our models together with the training mixtures are implemented in JAX/Flax (Bradbury et al., 2018) using the open-source T5X, SeqIO (Roberts et al., 2023) and Flaxformer (Heek et al., 2023) frameworks. For training, we use the Adafactor optimizer (Shazeer & Stern, 2018) with $\beta_{1}=0$, second order moment of $0.8$, and a language-model–style teacher forcing with softmax cross-entropy loss. For the learning rate schedule, we use the linear decay scheduler with a peak learning rate of $0.01$, 5k steps of warmup and a linear decay with decay factor of $63e-$ 6, and a dropout of $0.1$ on both ViT encoder and the mT5 encoder-decoder. We train our models for 340k steps with batch size $512$. With frozen ViT encoders, the training of a 340M parameter model (such as Small-i or Small-p) takes $\sim$33h on 64 TPU v5e chips and the training of a 1B parameter Large-i model takes $\sim$105h on 64 TPU v5e chips, both run on an internal TPU v5e cluster and should be reproducible with public GCP TPU v5e. Inference with a 340M parameter model takes around 1 hour for all 3 datasets, and 3 hours with Large-i.
Compute.
For the results reported in Section 4.6, we train one 340M parameter model per ablation. For entries with estimated variances, we train each model with 3 different seeds, respectively. This totals in $\sim$330h on 64 TPU v5e chips of training. We additionally train Small-p and Large-i to obtain the results in Table 2, accounting for $\sim$138h on 64 TPU v5e chips. In total, all experiments in this paper require a total of $\sim$500h of compute for training the models and running inference.
Decoding.
@@
Inference Speed.
We benchmarked inference speed on three different GPU devices by measuring tokens processed per second over five runs. The Titan RTX achieved an average of 114.43 tokens/s, the T4 averaged 47.14 tokens/s, and the V100 averaged 139.14 tokens/s. All measurements were conducted using greedy decoding under consistent conditions with batch size 32 across devices.
Appendix H General Virtual Sketching Framework Performance
@@
Appendix I Online Handwriting Recognizer Details
@@
Parameter IAMOnDB Small-i IAM derendered Small-i IAMOnDB + IAM derendered Small-p IAM derendered Small-p IAMOnDB + IAM derendered Large-i IAM derendered Large-i IAMOnDB + IAM derendered Dropout 0.25 0.25 0.3 0.3 0.25 0.25 0.25 Learning rate 0.005 0.001 0.001 0.005 0.005 0.001 0.001
Appendix J Expansion on Failure Cases
While our system exhibits satisfactory performance across various tested inputs, we have identified specific failure patterns that warrant further attention.
Extreme Stroke Width/Style Variations
As observed in Figures 46 and 44, performance can degrade when processing samples with stroke widths or styles significantly outside the training distribution. While training with our data augmentation improves robustness, it does not fully cover extreme cases. Enhancing robustness to diverse and extreme stylistic variations such as through the inclusion of harder examples in the training mixture or the use of expert models, remains an area for future work.
Disproportionate Component Scaling
While scaling the vision (ViT) component enhances detail preservation (Figure 5, Section 4.5.1), we observed that semantic understanding for tasks requiring recognition can be negatively impacted if the text-ink (mT5) component is not scaled proportionally. Finding the optimal balance between visual feature extraction and sequence modeling capabilities during scaling is an important consideration.
Vision Head Training Trade-offs
As detailed in Section 4.6, unfreezing the ViT heads allows capturing finer image details but introduces training instability and increased confusion between text and background noise. Future work could explore more sophisticated fine-tuning strategies, potentially adapting techniques like (Zhai et al., 2022) to balance detail capture with stability and semantic accuracy.
Untrained Languages
When presented with handwriting in a language completely absent from its training data (as illustrated with the Korean example for Small-p in Figure 40), the model cannot leverage semantic understanding. In these scenarios, it defaults to derendering based primarily on visual features and learned writing dynamics, which can lead to less accurate stroke reconstruction and an inability to utilize text-conditioned generation effectively. This highlights a limitation in applying text-conditioned modes to truly unseen scripts without further adaptation or fine-tuning.
Task Confusion during Decoding
Our multi-task training setup, while crucial for injecting priors, can occasionally lead to task confusion during decoding. We observed instances, particularly with the Recognize and Derender task, where the model is prompted to perform semantic understanding might mistakenly prioritize the recognition sub-task and fail to generate the corresponding ink sequence (we use greedy decoding), resulting in an empty ink output. This suggests potential ambiguity in the model’s internal representation or decoding process when asked to perform both tasks simultaneously. To mitigate this during inference, we implemented a fallback mechanism: if the Recognize and Derender task produces an empty ink output, we automatically re-run inference using the Vanilla Derender mode to ensure an ink sequence is generated based solely on the visual input. While effective as a practical workaround, addressing the root cause of this task confusion within the model architecture or decoding strategy represents an area for future investigation.
Appendix K Expansion on Out-of-Domain Generalization
In addition to the Generalization to Sketches as we described for Figure 7, we analyze more evidence of Out-of-Domain Generalization that we see in selected samples.
Generalization to untrained scripts and languages
As noted in Appendix J, even though the model was trained only on Chinese, English, and Vietnamese, it can still partially derender Korean text. A similar extension to German is shown in Figure 34, though in both cases the derendering quality is noticeably worse—especially when comparing Small-p with Small-i/Large-i. Because Small-i/Large-i was trained on both languages while Small-p was not, these examples highlight the model’s limited generalization ability to untrained scripts.
Generalization to unseen textures and backgrounds
We present additional evidence for generalization on a variety of textures and backgrounds, such as a neon sign on a window ledge in Figure 31 or handwriting written in sand in Figure 34.
Appendix L Human Evaluation
@@
Task setup You will be looking at input images of photos of handwriting in a variety of styles and the corresponding digital ink tracing performed either by a human or by one of our derendering models. Each stroke is rendered in a different color and with color gradient (darker $\to$ whiter) to reflect the direction. You will see the derendered ink as well as the derendered ink overlaid on top of the image and can use either to make the judgment. For each sample answer the two questions on the right and click submit to proceed to the next sample.
Is Reasonable Is the digital ink a reasonable tracing of the input image? Yes, it’s a good tracing It’s an okay tracing, but has some small errors It’s a bad tracing, has some major artifacts One interpretation of the quality of the ink could be like this: If you were to use it in a note taking app, would you keep it as is: good tracing need to make some edits: okay tracing not use it at all: bad tracing If you need some examples to get a better feeling, you can use the ones below, however there is some ambiguity within the task since we don’t want to impose any strict rules but allow for a natural variety of what people find a good derendering (it is designed to be subjective).
Is Real Could this digital ink have been produced by a human? Yes No In this question go with your intuition, do you think the ink could have been traced by a [reasonable] human?
@@
Appendix M Model Card
@@
Model Architecture A multimodal sequence-to-sequence Transformer (Vaswani et al., 2017) model with the mT5 (Xue et al., 2021) encoder-decoder architecture. It takes text tokens and ViT (Dosovitskiy et al., 2021) dense image embeddings as inputs to an encoder and autoregressively predicts discrete text and ink tokens with a decoder.
Input(s) A pair of image and text.
Output(s) Generated digital ink.
@@
Application The model is for research prototype, and the public version Small-p is planned to be released and available for the public.
Known Caveats None.
@@
System Description This is a standalone model.
Upstream Dependencies None.
Downstream Dependencies None.
@@
Hardware & Software Hardware: TPU v5e.
@@
For implementation details please refer to Appendix G.
Compute Requirements Please refer to Appendix G.
@@
Training Datasets The ViT encoder (Dosovitskiy et al., 2021) is pretrained on ImageNet-21k (Deng et al., 2009), mT5 encoder and decoder are initialized from scratch. The entire model is trained on the mixture of publicly available datasets described in Appendix C.
@@
Evaluation Methods Human evaluation (reported in Section 4.5.1) and automated evaluations (reported in Section 4.5.2).
@@
Sensitive Use The model is capable of converting images to digital inks. This model should not be used for any of the privacy intruding use cases, e.g. forging handwritings.
Known Limitations Reported in Section 4.7.
Ethical Considerations & Potential Societal Consequences Reported in Appendix O.
Appendix N Additional Visualizations
@@
N.1 More Visualizations on Public Evaluation Datasets
@@
Appendix O Impact Statements
We acknowledge the importance of considering the ethical implications of InkSight. This section outlines potential concerns and benefits to promote responsible use.
O.1 Ethical Considerations
Forgery
A key concern is forgery. However, InkSight is designed for offline-to-online handwriting conversion, not open-ended generation. We verify that proposed models are unable to perform open-ended ink generation. We provide random texts in task prompts of task Derender with Text (in Figure 3) with input images either empty or containing handwritings that do not match the label. We observe that the generated inks do not match the labels for all 100 cases inspected, signifying that the model is unable to generate the prompted text that is not present in the input image. More details in Section O.3.
Bias
Handwriting varies significantly across demographics, geographic regions, languages/scripts, age groups, and individuals with disabilities. Models trained primarily on data from dominant groups may exhibit bias, performing poorly for underrepresented styles or scripts. Ongoing efforts should focus on curating more diverse and representative training datasets encompassing a wider range of scripts, writing styles, and demographic groups. Evaluating model performance across different subgroups is crucial to identify and mitigate biases. Designing models with fairness metrics in mind is an important direction for future research.
O.2 Positive Societal Impacts
Despite the risks, InkSight offers significant potential benefits:
Document Digitization
Facilitates the preservation and indexing of personal notes, historical documents, and cultural heritage artifacts beyond their textual form.
Support for Low-Resource Languages
By enabling the conversion of existing handwritten documents in low-resource languages into digital ink, InkSight can help generate valuable training data for online handwriting recognition systems, improving digital tool availability for these communities.
Educational Tools
Digital ink conversion can potentially enable new interactive learning and teaching experiences for students and teachers.
Unsupervised Analysis via Geometric Pseudo-Labels
Beyond supervised training data generation, InkSight’s output offers a unique advantage for unsupervised analysis, particularly in low-resource or challenging scenarios. The generated digital ink sequence itself, even when semantic recognition is imperfect (e.g., for unseen languages or styles, as illustrated by the Korean example in Figure 40)captures valuable geometric and dynamic properties of the handwriting (shape, curvature, stroke connections). This ink sequence can serve as a powerful geometric pseudo-label. Offline handwriting images can then be clustered or compared based on the similarity of these ink pseudo-labels (e.g., using sequence comparison methods like DTW), leveraging purely structural information. This provides a language-agnostic alternative to OCR-based analysis, which relies on accurate text recognition and language models, allowing for grouping or analysis based on visual writing patterns even when the text content is unknown or poorly recognized.
O.3 Open-ended Generation
As illustrated in Figure 57, our experiments with Small-p, Small-i, and Large-i models demonstrate that when prompted with text not present in the original image, the models fail to generate accurate corresponding handwritten strokes. We regard this behavior as a natural safeguard against potential misuse, such as forgery or unauthorized reproduction of handwriting styles. Since the models can only reconstruct handwriting patterns from the input image rather than generating arbitrary new content, they inherently restrict attempts to synthesize fraudulent handwritten content.